If you used the doppler effect as a reason not to get a ticket for running a red light in the situation in no. 1, how much of a fine would you get for speeding? Assume that the fine is P50.00 for each kph (km per hour) that your speed exceeds the posted limit of 95 kph.


1
Expert's answer
2021-07-12T12:13:04-0400
"u=c\frac{f^2-f_0^2}{f^2+f_0^2}\\f=\frac{3\cdot10^8}{575\cdot10^{-9}}=5.21\cdot10^{14}Hz\\
f_0=\frac{3\cdot10^8}{700\cdot10^{-9}}=4.29\cdot10^{14}Hz\\""u=c\frac{5.21^2-4.29^2}{5.21^2+4.29^2}=0.192c\\"